(13) Th e probability density function of the loss X is /r(x) = { 0.02 (1-100), 0<x <100 100 0< otherwise The amount paid Y is 70 percent of that portion of the loss that e xceeds a deductible of 20. Determine E(Y). A)12 B)15 C)21。23 E) 25 (14) You are given that X is a normal randou variable such that E(X2) -1 A) 0.25 B) 0.33C) 0.64 (D)0.97 E) Does not exist P(X > 0) = 0.92. Find the variance 0.64 D)0.97 E) Does not exist the circled answers are wrong can show a good work on those two problems to arrive at correct answers please, will rate more points.
Answer #1

here E(Y)=0.7*(X-20)

therefore E(Y)=\int_{20}^{100} 0.7*(X-20)*f(x) dx =\int_{20}^{100}0.7*0.02*(x-20)*(1-x/100) dx

=\int_{20}^{100}0.014*(x-x2/100-20+x/5) dx =0.014*(x2/2-x3/300-20x+x2/10)|10020 =11.95~ 12

option A

14)

P(X>0)=0.92

for above z =-1.405

therefore \mu-1.405\sigma =0

\mu =1.405\sigma ..........(1)

E(X2) =\mu2+\sigma2 =1 .......(2)

solving above equations:

\sigma2 =0.33

option B